OpenStudy (anonymous):

The equation, y = 12841(1.01)x , represents the average cost of an education at a private college, after 1990. What is the average cost in 1990 (x = 0)? $12,969 $12,841 $13,122 $13,099

OpenStudy (anonymous):

is it 12,969? thats what i think it is

OpenStudy (campbell_st):

I think your equation should have a power of x... 12841(1.01)^x

OpenStudy (anonymous):

so 12841 you think?

OpenStudy (campbell_st):

well otherwise the cost in 1990 was $0 using your formula...

OpenStudy (anonymous):

1) Anything to the zero power is 1: y = 12841(1.01)^x y = 12841(1.01)^0 y = 12841(1) = $12841